x-3y-7z+w=11
2x-4y+6z-3w=-3
3x+2y+z+2w=19
4x+y-3z+w=22

Answer by Alan3354(69443)

x-3y-7z+w=11
2x-4y+6z-3w=-3
3x+2y+z+2w=19
4x+y-3z+w=22
---------------
w = 1207/373
x = 1641/373
y = -45/373
z = -160/373